OpenStudy (fellowroot):

given limit n goes to infinity of (Sn+1)/Sn = 1 >1 prove Sn goes to +infinity hint: limit as n goes to infinity of x^n = +infinity

OpenStudy (fellowroot):

its like a proof problem.

OpenStudy (anonymous):

Is \(S_n\) a sequence or a series?

OpenStudy (fellowroot):

sequence

OpenStudy (fellowroot):

\[\lim S _{n+1}/S _{n}\]

OpenStudy (zarkon):

why do you have 1>1

OpenStudy (fellowroot):

its L>1

OpenStudy (fellowroot):

if you want a pic i can post one

OpenStudy (zarkon):

I'll get you started. for large n he have \[s_{n+1}/s_n\approx L\] then \[s_{n+1}\approx L\cdot s_n\] \[s_{n+2}\approx L\cdot s_{n+1}\approx L^2\cdot s_{n}\] ... \[s_{n+k}\approx L^ks_{n}\] this should give you an idea of how to do this problem (in a formal way)

OpenStudy (fellowroot):

thanks i will use this and see if i can do it